# CodeAid: Evaluating a Classroom Deployment of an LLM-based Programming Assistant that Balances Student and Educator Needs

### 20 Jan 2024 v1

@@ 6.2.7. Help Write Code
Addressing Incorrect Responses Students typically reacted to CodeAid’s incorrect answers by asking follow-up questions and to provide clarification or more context. However, without the option to ask follow-up questions in the Help Fix Code feature, students sometimes entered a trial-and-error phase, repeatedly asking new questions with the same or slightly modified prompts. This highlights the need for intuitive interfaces in AI model control.
@@ 8. Beyond CodeAid: Implications for Pedagogical LLM-Powered Coding Assistants
The iterative development of CodeAid and insights gained from its semester-long deployment enable us to propose design implications for the broader design of AI assistants in educational contexts. We position these implications within four main stages of a student’s help seeking process with an AI assistant: 1) The decision to use the AI tool; 2) The formulation of a query; 3) The nature of response that is supplied; and 4) Actions needed once a response is received. Our results point to four high-level design considerations, for each of these stages, each with unique usability and educational trade-offs:
@@ 8.1. D1: Exploiting Unique Advantages of AI
An initial stage of the help-seeking process is deciding what learning resource to use within the learning material ecosystem that may be available to them. This leads to the first major design consideration of an AI coding assistant (Figure 7, D1): determining the role, scope, and unique advantages of future educational AI assistants in relation to other educational resources, like TA office hours, discussion boards, text books, etc. In our study, students appreciated several features unique to AI assistants, such as the ability to interact with the tool in natural language, the tool’s ability to provide contextual assistance, and tailored responses. More specific to CodeAid, student’s pinpointed unique advantages such as its stimulation of critical thinking (using pseudo-code), technical focus on C programming (relevant to their course), and its capability to “pull“ comprehensive documentation related to queries. Similar features should be built within future tools.
@@ 8.4. D4: Supporting Trust, Transparency and Control
This points to a fourth design consideration (Figure 7, D4): given the indirect nature of responses in educational AI assistants, how can we maximize user-experience, efficiency, and predictability of obtaining a helpful response. This could involve displaying the AI’s assumptions on input queries, engaging students in verifying them, and enabling feedback mechanisms to ensure high-quality responses. While such techniques could lead to higher accuracy, they may complicate user interaction and potentially overwhelm students. The best method for validating responses after generation is an open question, and requires targeted future work.
@@ 9. Limitations and Future Work
[…]
@@ 10. Conclusion
This paper presents the iterative design of CodeAid, an AI-powered coding assistant equipped with guardrails to prevent it from generating direct solutions in response to student queries. Instead, CodeAid provides scaffoldings, such as interactive pseudo-code, to foster engaging learning experiences. To understand how students utilize an AI-powered tutor and the broader implications of AI in scaling instructional expertise, we conducted a semester-long deployment of CodeAid in a programming class with 700 students. Our data collection included: (i) approximately 8,000 usages, coupled with students’ feedback on the responses’ usefulness; (ii) a thematic analysis of 1,749 usages in terms of correctness, helpfulness, and the extent of revealing direct solutions; (iii) weekly surveys and 22 semi-structured interviews with students; and (iv) a final anonymous survey focusing on the use of ChatGPT. Additionally, we interviewed eight programming educators to gather insights on the future of AI-powered educational tools. By synthesizing results from these diverse sources, we identified four high-level design considerations, including key design questions and trade-offs, in the emerging design space of educational AI tools. It is our hope the the results from our study, along with the broader design considerations which we have discussed, will help guide the future development of AI-powered codding assistants.

### 25 Feb 2024 v2

@@ 8. Beyond CodeAid: Implications for Pedagogical LLM-Powered Coding Assistants
The iterative development of CodeAid and insights gained from its semester-long deployment enable us to propose design implications for the broader design of AI assistants in educational contexts. We position these implications within four main stages of a student’s help-seeking process with an AI assistant: 1) The decision to use the AI tool; 2) The formulation of a query; 3) The nature of response that is supplied; and 4) Actions needed once a response is received. Our results point to four high-level design considerations, for each of these stages, each with unique usability and educational trade-offs:
@@ 8.1. D1: Exploiting Unique Advantages of AI
An initial stage of the help-seeking process is deciding what learning resource to use within the learning material ecosystem that may be available to them. This leads to the first major design consideration of an AI coding assistant (Figure 7, D1): determining the role, scope, and unique advantages of future educational AI assistants in relation to other educational resources, like TA office hours, discussion boards, textbooks, etc. In our study, students appreciated several features unique to AI assistants, such as the ability to interact with the tool in natural language, the tool’s ability to provide contextual assistance, and tailored responses. More specific to CodeAid, students pinpointed unique advantages such as its stimulation of critical thinking (using pseudo-code), technical focus on C programming (relevant to their course), and its capability to ”pull” comprehensive documentation related to queries. Similar features should be built within future tools.
@@ 8.4. D4: Supporting Trust, Transparency and Control
This points to a fourth design consideration (Figure 7, D4): given the indirect nature of responses in educational AI assistants, how can we maximize user experience, efficiency, and predictability of obtaining a helpful response. This could involve displaying the AI’s assumptions on input queries, engaging students in verifying them, and enabling feedback mechanisms to ensure high-quality responses. While such techniques could lead to higher accuracy, they may complicate user interaction and potentially overwhelm students. The best method for validating responses after generation is an open question, and requires targeted future work.
@@ 9. Limitations and Future Work
Our findings from the deployment of CodeAid are contextualized within a second-year C programming course at a single university, which may not directly generalize to other courses and contexts. Furthermore, the perceived accuracy and utility of CodeAid, which in turn influenced students’ trust and engagement, were heavily tied to the OpenAI models we used. The initial release of CodeAid with the [Uncaptioned image] model from 2021 could have adversely shaped students’ perceptions, particularly when compared to the subsequently adopted [Uncaptioned image] model from 2023. This factor might account for some of CodeAid’s inaccurate responses and the decline in student usage we observed over time. Additionally, the performance of these models varies between different programming languages (Xu et al., 2022), or even different code-related tasks (e.g. code explanation, fixing, or generation). For future work, we plan to run longitudinal studies to investigate how using pedagogical AI coding assistants affects long-term learning outcomes, competency, self-regulation abilities, and frustration levels in educational environments. More controlled studies could also be performed to directly compare these learning outcomes to the use of unrestricted LLMs like ChatGPT.
@@ 10. Conclusion
This paper presents the iterative design of CodeAid, an AI-powered coding assistant equipped with guardrails to prevent it from generating direct solutions in response to student queries. Instead, CodeAid provides scaffoldings, such as interactive pseudo-code, to foster engaging learning experiences. To understand how students utilize an AI-powered tutor and the broader implications of AI in scaling instructional expertise, we conducted a semester-long deployment of CodeAid in a programming class with 700 students. Our data collection included: (i) approximately 8,000 usages, coupled with students’ feedback on the responses’ usefulness; (ii) a thematic analysis of 1,749 usages in terms of correctness, helpfulness, and the extent of revealing direct solutions; (iii) weekly surveys and 22 semi-structured interviews with students; and (iv) a final anonymous survey focusing on the use of ChatGPT. Additionally, we interviewed eight programming educators to gather insights on the future of AI-powered educational tools. By synthesizing results from these diverse sources, we identified four high-level design considerations, including key trade-offs, in the emerging design space of educational AI tools. It is our hope the the results from our study, along with the broader design considerations which we have discussed, will help guide the future development of AI-powered codding assistants.